a car takes off from a stop sign with zero initial velocity but with nonzero acceleration of 6.50 m/s2.
a. what is the car's speed after 5.00 seconds?
b. What is the car's position after 5.00 seconds?
c. After the initial 5.00 seconds the car travels at a constant speed you found in (a). How long does it take for the car to travel 250 meters?
1
Expert's answer
2019-11-07T09:09:15-0500

(a) The speed of the car after 5 s


"v_f=v_i+at=0+6.50\:\rm m/s^2\times 5.0\: s=32.5\: m/s"

(b) The position of the car after 5 s


"x_f=x_i+v_it+\frac{at^2}{2}=\frac{6.50\:\rm m/s^2\times (5.0\:\rm s)^2}{2}=81.25\:\rm m"

(c) The time needed to travel 250 m with velocity 32.5 m/s


"t=\frac{d}{v}=\frac{250\:\rm m}{32.5 \:\rm m/s}=7.7\:\rm s"